# One Model to Train them All: Hierarchical Self-Distillation for Enhanced Early Layer Embeddings

### v1

@@ 2.4 Fine-tuning
To enhance representation learning, we replace the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36). We used a batch of 2 048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
@@
We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings. We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged. After conducting a grid search, we selected $51.0e-$ 5 as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
@@ 2.5 Evaluation
We evaluated ModularStarEncoder fine-tuned, on both text-to-code and code-to-code retrieval tasks using CodeXGLUE (Lu et al., 2021), which comprises several benchmarking datasets. For text-to-code retrieval, we employed the CodeSearchNet dataset, where the goal is to retrieve the most relevant code snippet given a natural language query. Specifically, the query corresponds to a documentation comment, and the model is tasked with ranking the correct code snippet among 999 distractor snippets Husain et al. (2019). This setup assesses the model’s ability to learn meaningful cross-modal representations between code and natural language.
For code-to-code retrieval, we relied on two datasets from CodeXGLUE: the Code Translation (CT) benchmark and POJ-104. The Code Translation dataset consists of semantically equivalent code snippets in different programming languages, and we framed the task as cross-language code retrieval rather than translation. In this setting, given a Java code snippet as a query, the model retrieves the corresponding C# implementation, testing its capability to capture cross-lingual semantic similarities between functionally equivalent programs.
In contrast, withPOJ-104 dataset, we want to evaluate the model on intra-language semantic search (POJ-104 contains only C++ snippets), where programs solve the same problem but with different implementations. This setup evaluates the model’s capacity to generalize across structural variations while preserving semantic equivalence.
Table 4: Performance of different models on Code Translation (CT) and POJ104 for code-to-code search with codeXGLUE dataset.
CT POJ104
Model MRR mAP
ModularStarEncoder 98.9 56.5
Codet5+ 110M-embedding 98.4 24.5
OpenAI text-embedding-3-large 98.8 82.9
Unixcoder 97.6 41.0
ModernBERT-large 93.1 27.3
3 Results and Discussion
@@
Table 5: Performance comparison of ModularStarEncoder layers and baseline fine-tuned models on the CodeSearchNet benchmark. The table displays the overall retrieval performance measured by Mean Reciprocal Rank (MRR). We refer to ModularStarEncoder, fine-tuned with multiple exit points simultaneously, as self-distilled. The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point.
@@
Model Size Ruby Javascript Go Python Java PHP avg. MRR
Layer-4 $\approx 160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
Layer-4 (self-distilled) 62.2 64.7 74.8 88.1 71.4 78.0 73.2
Layer-9 $\approx 300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
Layer-9 (self-distilled) 67.6 69.4 78.9 90.2 75.5 82.3 77.3
Layer-18 $\approx 550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
Layer-18 (self-distilled) 74.1 74.0 82.5 92.5 78.7 84.5 81.0
Layer-27 $\approx 800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
Layer-27 (self-distilled) 73.2 73.3 81.7 92.1 77.8 83.8 80.3
Layer-36 $\approx 1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
Layer-36 (self-distilled) 73.5 72.6 80.5 91.4 76.9 82.7 79.6
3.1 Benchmarks
Table 3 presents the results for CodeSearchNet (t2c) task in terms of Mean Reciprocal Rank (MRR) for each single language, average NDCG and average MRR. Results for Unixcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023). On CodeSearchNet, ModularStarEncoder achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (Wang et al., 2023) (770M), UniXcoder (Guo et al., 2022), and ModernBERT-large (Warner et al., 2024). The only encoder that surpasses ModularStarEncoder is OpenAI’s text-embedding-3-large.
Table 4 presents results from both POJ104 and CT datasets reported respectively in MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++# retrieval). ModularStarEncoder reaches the best performance among the tests. We decided to replicate the benchmarking for all models in a zero-shot setting for code-to-code tasks because our model does not integrate POJ104 and the code translation datasets in the training set.
Referring to Table 4, on the POJ104 dataset in zero-shot, ModularStarEncoder achieves an mAP of 0.57, which is state-of-art between open-sourced models, however it is significantly behind OpenAI text-embedding-3-large. We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
3.2 Ablation Study
We conducted an ablation study by fine-tuning singularly each exit point (also starting from ModularStarEncoder, pre-trained) and pruning the subsequent layers (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just one projection head on that layer). Finally, we compared the sliced models with the corresponding results (self-distilled) of the model fine-tuned with the multi-layer loss (ModularStarEncoder). ModularStarEncoder consistently outperforms the single-exit baseline, indicating that lower-level layers benefit from training signals propagated from deeper layers. This behavior is highlighted in Table 5, where ModularStarEncoder, indicated as self-distilled, outperforms all the single exit baselines consistently. This finding underscores a promising new direction in self-distillation for large-scale code and text models, enabling high performance even in more compact configurations. Moreover, Figure 4 illustrates that ModularStarEncoder maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval accuracy.
@@ 4 Related work
Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks. BERT (Feng et al., 2020) followed those findings, adapting the Transformer architecture (Vaswani et al., 2017) to enable a bi-directional representation with two different training objecting, namely the masked language modeling and the next sentence prediction losses. Lan et al. (2019); Liu et al. (2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization. More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
(a) MRR
(b) Recall@1
Figure 4: Performance Comparison Across Layers: The graph illustrates the MRR and the Recall@1 for different layers, comparing baseline models and a self-distilled model.
In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language. Feng et al. (2020) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT. This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al., 2020). GraphCodeBERT enhanced codeBERT (Feng et al., 2020) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection. UniXcoder (Guo et al., 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information. Wang et al. (2023) expanded these findings with codet5plus, stressing how multiple losses that leverage code semantics impact the model pertaining. The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation.
@@
Our work differs from previous work by adapting a modern architecture Lozhkov et al. (2024) to a code encoder-only based model and introducing a novel ’self-distillation’ mechanism. We replace the next sentence prediction loss with an in-context classification focused on the repository level and expand the context to 2048 tokens. Our novel self-distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
@@ 5 Conclusion
In this work, we introduced ModularStarEncoder, a modular multi-exit encoder architecture designed to improve efficiency and scalability in code retrieval tasks. By integrating an intra-model self-distillation mechanism, our approach enables multiple resolution models to be trained within a unified layer stack, reducing redundancy while maintaining high retrieval performance. Our evaluation on CodeXGLUE demonstrates that ModularStarEncoder achieves state-of-the-art results among open-source models, outperforming prior baselines across text-to-code and code-to-code retrieval tasks. Ablations further highlighted the benefits of self-distillation, showing that lower layers gain representational strength from deeper layers, leading to superior performance compared to single-exit models.
Beyond performance gains, ModularStarEncoder offers practical benefits by providing multiple exit points, allowing users to balance computational efficiency and accuracy based on resource constraints. The results suggest that self-distillation provides a promising direction for efficient large-scale encoders, reducing deployment costs without sacrificing effectiveness.
Finally, released in open-access our SynthCode2Code2NL and both pre-trained and fine-tuned ModularStarEncoder models.
@@ Acknowledgments
Limitations
Due to our dependence on multiple GPUs, we encountered significant computational constraints. Parameter grid searches with smaller and embryonic models were the only ways to extrapolate the best hyperparameter setup. The best hyperparameters for smaller models can differ from those for larger ones; thus, we faced a limitation in finding an optimal training setup. Ablating both the in-context classification and the multi-layer loss in a real scenario was impossible as we depended on smaller models to understand their performances. Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor undermines the possibility of replicating the experiments.
@@
When synthesizing the SynthCode2Code2NL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans. We tested the model’s performance on standard benchmarks. However, the impact of utilizing code snippets as synthetic data in training large language models for generalization over human text-to-code and code-to-code search is still not fully understood.
@@
Appendix A Synthetic dataset

### v2

@@ 4.1 Finetuning for Retrieval, text-to-code and code-to-code search
Using the SynthCoNL dataset detailed in section 5, we leveraged a instruction prompting approach Su et al. (2023) and trained a single model for both text-to-code and code-to-code retrieval tasks. The optimization objective uses CLIP loss (Radford et al., 2021) with the same multi-layer approach discussed in section 2.2, modified to enhance representation learning. Accordingly, we replaced the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36). We used a batch of 2,048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
@@
We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings. We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged. After conducting a grid search, we selected $1\mathrm{e}{-5}$ as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
Table 2: BigCloneBench code clone detection performance in a classification setup. MoSE maintains good performance (%) throughout different layers (L) with results on par with the state-of-the-art.
BigCloneBench
Model Recall Precision F1
MoSE (L4) 96.4 89.8 93.0
MoSE (L9) 96.6 90.4 93.4
MoSE (L18) 96.4 92.1 94.2
MoSE (L27) 96.5 91.8 94.1
MoSE (L36) 96.5 91.7 94.1
UniXcoder 92.9 97.6 95.2
CodeBERT 94.7 93.4 94.1
CodeT5+ (770M) 96.7 93.5 95.1
4.2 Finetuning for classification, code clone detection
[…]
5 SynthCoNL
Prompt Translate this “‘print("Hello World")”’ from Python to Rust. Here is the translated code: Qwen2.5-Coder-7B-Instruct Output fn main() { println!("Hello World!"); }
We built SynthCoNL, a dataset that supports training and evaluation of encoders in text-to-code and code-to-code search, under multiple languages. Using the popular CodeSearchNet (Husain et al., 2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript. We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig. 3). During code translation, we carried out a greedy sampling so as to prevent semantic discrepancies.
This process yielded pairs of code snippets in distinct languages tied to the same natural language description. As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages. SynthCoNL contains 1,071,367 triplets where, in the first code column, we directly sampled code snippets from CodeSearchNet, including Python, Java, PHP, and Go. The second code column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets. After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments. Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port). During the preprocessing phase of SynthCoNL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column. During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams.
6 Results and Discussion
@@ 6.1 Evaluation
Figure 4: Performance of different models on CT and POJ104 for code-to-code retrieval with CodeXGLUE dataset. MoSE has state-of-the-art results with the CT benchmark and outperforms all the open-source models over the POJ104 benchmark. Compared to closed-source solutions (OpenAItext-embedding-3-large), MoSE performs on par in the code translation task and strongly reduces the gap between open- and closed-source models with the POJ104 benchmark.
We evaluated MoSE, after fine-tuned for retrieval, on diverse code understanding tasks, leveraging the CodeXGLUE (Lu et al., 2021) group of benchmarks.
On CodeSearchNet dataset, we assessed cross-modal (natural language to code) learning by retrieving target code from 999 distractors using natural language queries. For code-to-code retrieval, again against 999 distractors, we utilized CodeXGLUE’s Code Translation (CT) dataset to test cross-lingual retrieval (e.g., Java to C# snippets implementing the same functionality), and its POJ104 dataset (C++ snippets, semantically equivalent but syntactically different) for intra-lingual semantic search, evaluating generalization over structural differences while preserving semantics. In our retrieval performance evaluation, when comparing MoSE with other encoders, we included results from OpenAI’s text-embedding-3-large as a representative commercial, closed-source model.
Additionally, MoSE was fine-tuned and evaluated on the separate BigCloneBench (Svajlenko & Roy, 2015) dataset for Java code clone detection (a binary classification task), assessing its classification performance.
6.2 Benchmarks
Table 1 presents the results for the CodeSearchNet (text-to-code) task in terms of Mean Reciprocal Rank (MRR, higher is better) for each single language, average Normalized Discounted Cumulative Gain (NDCG, higher is better), and average MRR. Results for UniXcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023). On CodeSearchNet, MoSE achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (770M), UniXcoder, and ModernBERT-large. The only encoder that surpasses MoSE is OpenAI’s text-embedding-3-large.
In Figure 4, we present results from both POJ104 and CT datasets reported respectively in terms of MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++ retrieval). MoSE reaches the best performance among the evaluated models. We further replicated the benchmarking for all models in a zero-shot setting for code-to-code tasks as our model does not integrate POJ104 and the code translation datasets in the training set. We present a comprehensive and detailed analysis of code retrieval errors in appendix C.
Figure 5: Similarity score Heatmap presenting results from permutation tests (10,000 permutations with $\alpha$ = 0.05) across different exit points for Python-to-Java retrieval. * indicates p < 0.05 and ** indicates p < 0.001. Different layers, despite a common training objective, yield different similarity scores.
Referring to Figure 4, on the POJ104 dataset in zero-shot, MoSE achieves a mean Average Precision (mAP, higher is better) of 75.9, which is state-of-the-art among open-sourced models. However, it is significantly behind OpenAI text-embedding-3-large. We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
We observe a marked difference in performance depending on whether the task is framed as text-to-code or code-to-code retrieval. As shown in Figure 1(b), shallow layers yield better performance in the code-to-code setting, whereas for text-to-code retrieval, optimal performance emerges at layer 18.
This representation divergence suggests that the input modality influences where semantic alignment is most effectively captured within the model. This observation aligns with Valeriani et al. (2023) findings, underlining that semantic information is distributed unevenly across layers and is better expressed in the intermediate ones. MoSE leverages this property by supporting dynamic exit points, enabling users to balance computational cost and task performance based on where the most relevant representations emerge.
To further investigate this phenomenon, we compared similarity scores across layers to assess how each layer represents code snippets differently. Figure 5 shows permutation test results for a Python-to-Java retrieval task, highlighting significant variations in similarity scores that support our previous findings. A detailed analysis of these results is provided in appendix C.
In Table 2, we report classification performance across different exit layers of MoSE for the BigCloneBench benchmark. Results show stable performance throughout layers, with F1-scores from Layer 18 onward (94.2–94.1) closely matching those of deeper exits. Earlier layers (e.g., Layers 4 and 9) exhibit slightly lower precision, reflecting the expected trade-off between early inference and representational depth.
Our method remains competitive compared to state-of-the-art models. While CodeT5+ (770M) and UniXcoder achieve peaks F1 of 95.1 and 95.2, MoSE delivers on-par results (-0.9) by Layer 18, with the added benefit of a modular architecture.
6.3 Ablation Study
We conducted an ablation study for the Multi-Layer loss, by fine-tuning (starting from MoSE, pre-trained) on SynthCoNL, each exit point singularly (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just the projection head on that layer). Finally, we compared the baseline models with the corresponding results (Self-Distilled) of MoSE fine-tuned for retrieval with the multi-layer loss; all tests are conducted on the CodeSearchNet dataset. MoSE consistently outperforms the single-exit baselines (up to +4.36% averaged Recall@1 gain on layer 9), indicating that lower-level layers benefit from training signals propagated from deeper layers. This behavior is highlighted in Table 3, where MoSE, indicated as Self-Distilled, outperforms all the single exit baselines consistently. This finding underscores a promising new direction in Self-Distillation for large-scale code and text models, enabling high performance even in more compact configurations. Moreover, Figure 6 illustrates that MoSE maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval performances.
@@ 6.3 Ablation Study
Table 3: Performance comparison of MoSE layers and baseline fine-tuned models on the CodeSearchNet benchmark. The table displays the overall retrieval performance measured by the Mean Reciprocal Rank (MRR, higher is better). We refer to MoSE, fine-tuned with multiple exit points simultaneously, as Self-Distilled. The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point. MoSE fine-tuned with multi-layer loss (Self-Distillation) consistently outperforms the baselines, showing a more significant difference in the initial layers.
@@
Model type Num. Layers Num. Params. Ruby Javascript Go Python Java PHP avg. MRR
Base 4 $160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
Self-Distilled 62.2 64.7 74.8 88.1 71.4 78.0 73.2
Base 9 $300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
Self-Distilled 67.6 69.4 78.9 90.2 75.5 82.3 77.3
Base 18 $550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
Self-Distilled 74.1 74.0 82.5 92.5 78.7 84.5 81.0
Base 27 $800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
Self-Distilled 73.2 73.3 81.7 92.1 77.8 83.8 80.3
Base 36 $1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
Self-Distilled 73.5 72.6 80.5 91.4 76.9 82.7 79.6
@@ 7 Related work
Figure 6: Performance Comparison between MoSE fine-tuned with Multi-layer loss (Self-Distilled model) and baselines fine-tuned just on single exit points: The graph illustrates averaged MRR and Recall@1 results for different layers over the CodeSearchNet test set. Layers fine-tuned with Multi-Layer loss outperforms the baselines (up to 3.47% MRR and +4.36% Recall@1).
Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks. BERT (Feng et al., 2020) followed those findings, adapting the Transformer architecture (Vaswani et al., 2017) to enable a bi-directional representation with two different training objecting, namely the masked language modeling and the next sentence prediction losses. Lan et al. (2019) and Liu et al. (2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization. More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language. Feng et al. (2020) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT. This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al., 2020). GraphCodeBERT enhanced codeBERT (Feng et al., 2020) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection. UniXcoder (Guo et al., 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information. Wang et al. (2023) expanded these findings with CodeT5+, stressing how multiple losses that leverage code semantics impact the model pertaining. The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation.
@@
Our work differs from previous work by adapting a modern architecture (StarCoder-2 (Lozhkov et al., 2024)) to a code encoder-only based model and introducing a novel ’Self-Distillation’ mechanism. We replace the next sentence prediction loss with an In-Context Classification focused on the repository level and expand the context to 2,048 tokens. Our novel Self-Distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
@@ 8 Conclusion
We introduced MoSE, a modular, multi-exit encoder that uses Self-Distillation to enhance early-layer representations while achieving state-of-the-art results on code understanding benchmarks. By integrating multiple exit points, MoSE enables explicit efficiency-accuracy trade-offs tailored to deployment constraints. Our approach not only allows for modular use but also enhances the alignment between task requirements and representation depth. This improves task-specific layer specialization, providing the end-user with a more adaptable model.
Future work will explore combining exit-points and further understanding task type and depth interactions. While our results are encouraging, we acknowledge that computational constraints limited extensive hyperparameter tuning and pre-training ablations, and that fine-tuning with synthetic code generated from code translation has an unclear impact. We discuss these limitations in detail in appendix A.
@@ Acknowledgments
This work was supported by Future AI Research (FAIR) PE01, SPOKE 8 on PERVASIVE AI funded by the National Recovery and Resilience Plan (NRRP).
Appendix A Limitations
Due to our dependence on multiple GPUs, we encountered significant computational constraints. Parameter grid searches using smaller and earlier models were the only methods to extrapolate the best hyperparameter setup. The optimal hyperparameters for smaller models can differ from those for larger ones; consequently, we encountered a limitation in discovering an effective training setup. Conducting an ablation study for the in-context classification loss was not practicable due to the high costs of the pre-training stage. Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor can also undermine the possibility of replicating the experiments.
@@ Appendix A Limitations
When synthesizing the SynthCoNL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans. We tested the model’s performance on standard benchmarks. However, the impact of utilizing code snippets as synthetic data in training large language models for generalization over human text-to-code and code-to-code search is still not fully understood.
@@
Appendix B Synthetic dataset
Table 4: SynthCoNL composition: Average character count and sample size per language for the CodeSearchNet seed dataset and synthesized portion obtained through translation.
Language CSN samples CSN avg. char Synth. samples Synth. avg. char
English 1,071,367 180 - -
PHP 280,706 514 116,967 579
Python 274,454 474 117,374 518
Go 234,089 350 124,125 541
Java 282,118 505 116,098 707
C++ - - 141,956 938
Ruby - - 158,494 456
C - - 136,365 1,029
JavaScript - - 159,988 557